# Vision-Language Models for Medical Report Generation and Visual Question Answering: A Review

### v1

@@ Abstract
Medical vision-language models (VLMs) combine computer vision and natural language processing to analyze visual and textual medical data. Our paper reviews recent advancements in developing VLMs specialized for healthcare, focusing on models designed for medical report generation and visual question answering. We provide background on natural language processing and computer vision, explaining how techniques from both fields are integrated into VLMs to enable learning from multimodal data. Key areas we address include the exploration of medical vision-language datasets, in-depth analyses of architectures and pre-training strategies employed in recent noteworthy medical VLMs, and comprehensive discussion on evaluation metrics for assessing VLMs’ performance in medical report generation and visual question answering. We also highlight current challenges and propose future directions, including enhancing clinical validity and addressing patient privacy concerns. Overall, our review summarizes recent progress in developing VLMs to harness multimodal medical data for improved healthcare applications.
@@ 1 Introduction
The last decade has witnessed enormous progress in artificial intelligence (AI) and machine learning (ML), including the development of foundation models (FMs), large language models (LLMs), and vision-language models (VLMs). These AI/ML developments have started transforming several aspects of our daily lives, including healthcare. AI/ML can potentially transform the whole continuum of healthcare by significantly optimizing and improving disease screening and diagnostic procedures, treatment planning, and post-treatment surveillance and care [Baj+21]. Various computer vision (CV) and natural language processing (NLP) models, more recently LLMs, have been instrumental in driving this transformative trend [He+23, Zho+23]. CV models have been trained and validated for various screening and diagnosis use cases leveraging radiology data from X-rays, mammograms, magnetic resonance imaging (MRI), computed tomography (CT), and others. Recently, AI/ML models focused on digital pathology using histopathology and immunohistochemistry data have also shown significant advances in accurate disease diagnosis, prognosis, and biomarker identification. On the other hand, by training models using large datasets of medical literature, clinical notes, and other healthcare-related text, LLMs are able to extract insights from electronic health records (EHR) efficiently, assist healthcare professionals in generating concise summary reports and facilitate the interpretation of patient information. Noteworthy examples of such LLMs include GatorTron [Yan+22], ChatDoctor [Li+23b], Med-PaLM (Medical Pathways Language Model) [Sin+23] and Med-Alpaca [Han+23].
@@
The healthcare data is inherently multimodal, and consequently, the AI/ML models often need to be trained using multiple data modalities, including text (e.g., clinical notes, radiology reports, pathology reports, etc.) imaging (e.g., radiology scans or histopathology slides), and tabular data (e.g., numerical data such as vitals or labs and categorical data such as race, gender, and others) [Aco+22, Shr+23, Waq+23, Tri+23]. In routine clinical practice, healthcare professionals utilize a combination of these data modalities for diagnosing and treating various conditions. Integrating information from diverse data modalities enhances the precision and thoroughness of disease assessments, diagnoses, treatment planning, and post-treatment surveillance. The need for AI/ML models to ingest, integrate, and learn from information stemming varied data sources is the driving force for multimodal learning [Hua+21].
@@
The recent progress in multimodal learning has been driven by the development of vision-language models (VLMs) [Gan+22, Che+23]. These cutting-edge models can analyze, interpret, and derive insights from both visual and textual data. In the medical domain, these models are contributing to developing a more holistic understanding of patient information and improving performance of ML models in various clinical tasks. Many of these models, like CLIP (Contrastive Language–Image Pre-training) [Rad+21], LLaVa (Large Language and Vision Assistant) [Liu+23a], and Flamingo [Ala+22] are tailored to healthcare domain through training on extensive medical datasets. Adapting VLMs for medical visual question-answering [Lin+23a] is particularly noteworthy, empowering healthcare professionals to pose queries regarding medical images such as CT scans, MRIs, X-rays, and more. The question-answering capability elevates the interactive nature of the AI/ML models in healthcare, facilitating dynamic and informative exchanges between healthcare providers and the AI system. Furthermore, adapting VLMs for medical report generation enables them to produce detailed and contextually relevant reports by amalgamating information from both visual and textual sources. This not only streamlines the documentation process but also ensures that the generated reports are comprehensive and accurately reflect the subtleties present in the data, further enhancing healthcare workflow efficiency.
1.1 Contribution & Organization
The contribution and organization of this review is as follows:
  1. 1.
    It offers a background on Neural Networks (NNs), Natural Language Processing (NLP), and Computer Vision (CV), ensuring accessibility for readers without a machine learning background (Section 2).
  2. 2.
    It delves into a comprehensive exploration of Vision-Language Models (VLMs), offering detailed insights into their adaptation for downstream tasks (Section 3).
  3. 3.
    It compiles the first and comprehensive list of diverse vision-language datasets used for medical report generation and visual question answering accompanied by detailed dataset descriptions (Section 4).
  4. 4.
    It compiles the first and detailed list of metrics that are used for VLM evaluation (Section 5).
  5. 5.
    It presents a comprehensive review of VLMs specifically tailored for medical report generation and visual question answering, which is the first of its kind to the best of our knowledge. It also includes a comparison of medical VLMs among each other (Section 6).
  6. 6.
    It discusses current challenges in the field and proposes various potential research directions that could significantly shape the future of medical VLMs (Section 7).
The overall structure of this review is shown in Fig. 1.
Figure 1: The organization of the review paper.
2 Basic Machine Learning Background
We start this section with a background on neural networks (NNs), natural language processing (NLP) and computer vision (CV).
2.1 Neural Networks
@@
Machine learning (ML) and artificial intelligence (AI), as we understand them today, began to take shape in the late 1940s and early 1950s [Bal21]. Neural networks (NNs) stand out as classical machine learning models, drawing inspiration from the structure and functioning of the human brain. They are composed of layers of interconnected nodes or neurons, arranged into an input layer, an output layer, and multiple intermediate layers called hidden layers. The basic NN is a feedforward NN, where neurons can be numbered in such a way that a connection from neuron $i$ to neuron $j$ can exist if and only if $i<j$ [Bal21]. In any NN, the connections between nodes carry weights, and neurons utilize activation functions on their inputs. Activation functions play a crucial role in introducing non-linearity to the model, enabling it to learn complex mappings between inputs and outputs. Common activation functions include the sigmoid, hyperbolic tangent (tanh), and Rectified Linear Unit (ReLU). NNs utilize a loss function to quantify the difference between the predicted output and the actual target values. The goal during training is to minimize this loss.
@@
Backpropagation, short for ”backward propagation of errors,” is a key algorithm for training deep neural networks. It involves two main steps: forward pass and backward pass. In the forward pass, input data is fed through the network, and predictions are generated. The backward pass calculates the gradient of the loss function with respect to the weights of the network. This gradient is then used to update the weights, minimizing the difference between predicted and actual values. Backpropagation is a straightforward application of the chain rule for computing derivatives [Bal21]. After the backward pass, the optimization algorithm takes these gradients and adjusts the weights and biases of the NN to minimize the loss. Among common optimization methods are stochastic gradient descent (SGD) [Rob51] and Adam (Adaptive Moment Estimation [KB14].
In practice, several issues can be encountered when training NNs. The vanishing gradient problem arises when gradients become extremely small as they are propagated backward through the layers during training. As a result, the weights in the early layers of the network receive negligible updates, hindering their learning. This phenomenon is particularly problematic for deep architectures with many layers. To tackle this issue, ReLU activation functions or skip connections, as proposed in ResNet [He+16], are commonly utilized. Another common issue is the exploding gradient problem. This occurs when gradients become excessively large during backpropagation, leading to unstable training and convergence issues. To address exploding gradients, techniques such as gradient clipping, which involves scaling gradients if they exceed a certain threshold, are employed. Moreover, overfitting and underfitting are common issues in NN training. Overfitting occurs when the model performs well on the training data but poorly on new, unseen data, indicating that it has memorized the training set rather than learning the underlying patterns. Increasing model complexity is a common strategy to address overfitting. On the other hand, underfitting arises when the model’s performance is poor on both the training and testing data, indicating a failure to capture the underlying patterns in the data. To mitigate these problems, regularization methods such as dropout, which randomly drops neurons during training to prevent reliance on specific features, are commonly employed [Bal21].
2.2 Natural Language Processing
@@
Natural Language Processing (NLP) is the analysis of linguistic data, most commonly in the form of textual data such as documents or publications, using computational methods [VC13]. NLP encompasses a variety of tasks aimed at understanding, processing, and generating human language. Named entity recognition (NER), a prominent NLP task, focuses on identifying and classifying entities within the text, such as names of individuals, medical conditions, etc. For instance, in medical literature, NER can assist in extracting crucial information from documents. Text summarization in NLP is widely used for generating coherent summaries of lengthy texts. Sentiment analysis is a task that determines the emotional tone expressed in a given text, providing valuable insights for applications like social media monitoring or customer feedback analysis. Machine translation is a fundamental NLP task in breaking down language barriers by automatically translating text from one language to another. Question answering in NLP is directed at comprehending and responding to user queries, propelling advancements in virtual assistants and information retrieval.
@@
The first step in NLP is tokenization, which is the mechanism of splitting or fragmenting the sentences and words to their possible smallest morpheme called a token. A morpheme is the smallest possible word after which it cannot be broken further [RB21]. One example of a word-level tokenization method is whitespace tokenization, which segments text based on whitespace characters. In many NLP applications, subword tokenization methods are preferred due to their effectiveness in handling out-of-vocabulary words. WordPiece [Wu+16] begins by treating each character as a token, creating an initial vocabulary. Employing a flexible and adaptive merging strategy, WordPiece considers any pair of adjacent characters or subword units that enhance the overall likelihood of the training data. This likelihood reflects the model’s probability of accurately representing the training data given its current state. In contrast, Byte Pair Encoding (BPE) [SHB16] shares similarities with WordPiece but adheres to a more deterministic merging strategy. In each iteration, BPE merges the most frequent pair of adjacent characters or subword units, progressing toward a predefined vocabulary size. Byte level BPE [WCG20] operates at an even finer granularity, considering individual bytes rather than characters. Byte Level BPE extends the concept of subword tokenization to bytes, allowing it to capture more nuanced patterns at the byte level.
@@
Tokens are then often transformed into numerical vectors that capture semantic relationships; this feature extraction is called a word embedding. Word embeddings are commonly used in tasks where understanding the meaning of words in context is crucial, for example, in sentiment analysis, named entity recognition, and language translation.Word2Vec [Mik+13] is a widely used word embedding technique that uses two models: Skip-Gram [Mik+13a] and Continuous Bag of Words (CBOW) [Mik+13b]. In Skip-Gram, the model predicts context words given a target word, capturing semantic associations. Conversely, CBOW predicts the target word based on its context, emphasizing syntactic structures. In both models, the “context word” refers to words within a specified window around the target word, and the “target word” is the word for which predictions are made. Word2Vec is computationally efficient, making it well-suited for large datasets and general-purpose applications. GloVe (Global Vectors) [PSM14] is a word embedding model that distinguishes itself by capturing global semantic relationships. It focuses on the entire corpus rather than local context windows. The model builds a co-occurrence matrix based on the global statistics of word pairs and then employs an objective function to generate word vectors that reflect the ratio of co-occurrence probabilities. GloVe uses an implicit Skip-Gram approach, capturing word relationships globally, making it ideal for tasks that demand a holistic understanding of word connections. FastText [Boj+17] is another word embedding that is particularly effective for handling out-of-vocabulary words and morphologically rich languages. It adopts a sub-word approach, breaking words into “n-grams”, and employs a Skip-Gram training method similar to Word2Vec [Mik+13] to learn embeddings for these sub-word units. There are also word embeddings tailored to better represent biomedical and clinical terms. In tasks where the order of words is not essential other feature extraction techniques can be effective, for example, bag-of-words (BoW) in text classification or term frequency-inverse document frequency (tf-idf) for information retrieval.
@@
In addition to general-purpose word embeddings, there are ones designed for biomedical and clinical terms. BioWordVec [Zha+19] incorporates MeSH (Medical Subject Headings) terms along with text from PubMed abstracts and employs the fastText [Boj+17] algorithm to learn improved biomedical word embeddings. Another prominent approach is Cui2vec [Bea+20], which leverages diverse multi-modal data derived from medical publications and clinical notes. Cui2vec systematically maps medical terms onto a common Concept Unique Identifier (CUI) space, followed by the construction of a co-occurrence matrix. This matrix captures instances where different CUIs appear together, which is a foundation for generating word embeddings using techniques such as GloVe [PSM14] or Word2Vec [Mik+13]. It’s worth noting that, post word embedding extraction, the incorporation of positional encodings becomes common to address the challenge of representing the order of elements in a sequence. These encodings, often based on sinusoidal functions, systematically encode token positions, enriching embeddings with positional information for utilization in machine learning models tailored to specific NLP tasks.
@@
Recurrent Neural Networks (RNNs) are widely employed for pattern detection in sequential data, encompassing diverse types such as genomic sequences, text, or numerical time series [Sch19]. Operating on the principle of preserving a form of memory, RNNs incorporate a cyclic structure by looping the output of a specific layer back to the input, facilitating the prediction of subsequent layer outputs. This mechanism empowers RNNs to adeptly model sequential and temporal dependencies, capturing information from preceding time steps within hidden states. Despite their capacity to retain information from past inputs, RNNs encounter challenges in preserving long-term dependencies within input sequences due to the vanishing gradient problem. To address this, several RNN variants, including Long Short-Term Memory (LSTM) [HS97] and Gated Recurrent Unit (GRU) [Cho+14] networks, have been designed to enhance their ability to capture and utilize long-range dependencies in sequential data.

### v2

# references.bib

@@ Abstract
Medical vision-language models (VLMs) combine computer vision (CV) and natural language processing (NLP) to analyze visual and textual medical data. Our paper reviews recent advancements in developing VLMs specialized for healthcare, focusing on models designed for medical report generation and visual question answering (VQA). We provide background on NLP and CV, explaining how techniques from both fields are integrated into VLMs to enable learning from multimodal data. Key areas we address include the exploration of medical vision-language datasets, in-depth analyses of architectures and pre-training strategies employed in recent noteworthy medical VLMs, and comprehensive discussion on evaluation metrics for assessing VLMs’ performance in medical report generation and VQA. We also highlight current challenges and propose future directions, including enhancing clinical validity and addressing patient privacy concerns. Overall, our review summarizes recent progress in developing VLMs to harness multimodal medical data for improved healthcare applications.
@@ 1 Introduction
The last decade has witnessed enormous progress in artificial intelligence (AI) and machine learning (ML), including the development of foundation models (FMs), large language models (LLMs), and vision-language models (VLMs). These AI/ML developments have started transforming several aspects of our daily lives, including healthcare. AI/ML can potentially transform the whole healthcare continuum by significantly optimizing and improving disease screening and diagnostic procedures, treatment planning, and post-treatment surveillance and care [bajwa2021aihealthcare]. Various computer vision (CV) and natural language processing (NLP) models, more recently LLMs, have been instrumental in driving this transformative trend [he2023llmsurvey, zhou2023llmsurvey2]. CV models have been trained and validated for various screening and diagnosis use cases leveraging radiology data from X-rays, mammograms, magnetic resonance imaging (MRI), computed tomography (CT), and others. Recently, AI models focused on digital pathology using histopathology and immunohistochemistry data have also shown significant advances in accurate disease diagnosis, prognosis, and biomarker identification [WAQAS2023100255]. On the other hand, by training models using large datasets of medical literature, clinical notes, and other healthcare-related text, LLMs can extract insights from electronic health records (EHR) efficiently, assist healthcare professionals in generating concise summary reports, and facilitate the interpretation of patient information. Noteworthy examples of such LLMs include GatorTron [yang2022gatortron], ChatDoctor [li2023chatdoctor], Med-PaLM (Medical Pathways Language Model) [singhal2023medpalm] and Med-Alpaca [han2023medalpaca].
@@
The healthcare data is inherently multimodal, and consequently, the AI/ML models often need to be trained using multiple data modalities, including text (e.g., clinical notes, radiology reports, surgical pathology reports, etc.), imaging (e.g., radiological scans, digitized histopathology slides, etc.), and tabular data (e.g., numerical data such as vitals or labs and categorical data such as race, gender, and others) [acosta2022biomedicalai, shrestha2023medicalvlms, waqas2023multimodaloncology, tripathi2023multimodaloncologydataset, Mohsan2023]. In routine clinical practice, healthcare professionals utilize a combination of these data modalities for diagnosing and treating various conditions. Integrating information from diverse data modalities enhances the precision and thoroughness of disease assessments, diagnoses, treatment planning, and post-treatment surveillance. The need for AI/ML models to ingest, integrate, and learn from information stemming from varied data sources is the driving force for multimodal learning [huang2021multimod, waqas2023multimodaloncology].
@@
The recent progress in multimodal learning has been driven by the development of vision-language models (VLMs) [gan2022vlmsurvey, chen2022vlp, Mohsan2023]. These cutting-edge models can analyze, interpret, and derive insights from both visual and textual data. In the medical domain, these models contribute to developing a more holistic understanding of patient information and improving the performance of ML models in various clinical tasks. Many of these models, like CLIP (Contrastive Language–Image Pre-training) [radford2021clip], LLaVa (Large Language and Vision Assistant) [liu2023llava], and Flamingo [alayrac2022flamingo] are tailored to healthcare domain through training on extensive medical datasets. Adapting VLMs for medical visual question-answering [lin2023medvqa] is particularly noteworthy, empowering healthcare professionals to pose queries regarding medical images such as CT scans, MRIs, mammograms, ultrasound, X-rays, and more. The question-answering capability elevates the interactive nature of the AI/ML models in healthcare, facilitating dynamic and informative exchanges between healthcare providers and the AI system. Furthermore, adapting VLMs for medical report generation enables them to produce detailed and contextually relevant reports by amalgamating information from both visual and textual sources. This not only streamlines the documentation process but also ensures that the generated reports are comprehensive and accurately reflect the subtleties present in the data, further enhancing healthcare workflow efficiency.
In contrast to previous related surveys [lin2023medvqa, ting2023mrgsurvey, shrestha2023medicalvlms], this review focuses on the latest advancements in VLMs tailored for medical report generation and visual question-answering. The overall structure of this review is shown in Fig. 1 and is outlined as follows. In Section 2, we provide essential background on neural networks, CV, and NLP. In Section 3, we delve into the exploration of VLMs’ architectures, training strategies, and downstream tasks. The goal of Section 2 and Section 3 is to ensure the accessibility of this review for readers, irrespective of their ML background. We split Section 4 into three key sub-sections. In Section 4.1, we describe 17 publicly available vision-language datasets. These datasets encompass medical image-text pairs or question-answer pairs related to medical images. Next, in Section 4.2, we meticulously outline the metrics and their formulas, where applicable, employed for evaluating VLMs in the context of report generation and visual question-answering tasks. In Section 4.3, we conduct a thorough review of 15 recent medical VLMs, with 14 of them being publicly available. To the best of our knowledge, most of these models have not been reviewed in any previous surveys. Finally, in Section 5, we discuss the current challenges within the field of medical VLMs, offering insights into potential research directions that could profoundly influence their future development. The list of medical VLMs and datasets can also be found on GitHub.
Figure 1: Organization of the review paper.
2 Machine Learning (ML) - A Brief Review
2.1 Neural Networks (NNs)
@@
ML and AI, as we understand them today, began to take shape in the late 1940s and early 1950s [baldi2021dlscience]. NNs stand out as classical ML models, drawing inspiration from the structure and functioning of the human brain. They are composed of layers of interconnected nodes or neurons arranged into an input layer, an output layer, and multiple intermediate layers called hidden layers. The basic NN is a “feedforward NN”, where neurons can be numbered in such a way that a connection from neuron $i$ to neuron $j$ can exist if and only if $i<j$ [baldi2021dlscience]. In any NN, the connections between nodes carry weight, and neurons utilize “activation functions” on their inputs. Activation functions play a crucial role by introducing non-linearity to the model, enabling it to learn complex nonlinear mappings between inputs and outputs. Common activation functions include the sigmoid, hyperbolic tangent (tanh), and Rectified Linear Unit (ReLU). NNs utilize a loss function to quantify the difference between the predicted outputs and the actual targets. The loss function produces a scalar value, and the goal during training is to minimize this loss value.
@@
Backpropagation, short for backward propagation of errors, is a key algorithm for training deep neural networks. During the forward pass, input data is fed through the network, predictions are generated, and a scalar loss value is calculated using the loss function. During backpropagation, we calculate the gradient of the loss function with respect to the weights of the network. This gradient information is then used to update the weights in an effort to minimize the difference between predicted and target values. Backpropagation is an application of the chain rule for computing derivatives [baldi2021dlscience]. After the backward pass, the optimization algorithm takes these gradients and adjusts the learnable parameters (weights and biases) of the NN, which, in turn, will result in the minimization of the loss value in the next batch. Common optimization methods include gradient descent, stochastic gradient descent (SGD), [robbins1951sgd], Adam (Adaptive Moment Estimation [kingma2014adam], and many others.
2.2 Natural Language Processing (NLP)
@@
NLP is the analysis of linguistic data, most commonly in the form of textual data such as documents or publications, using computational methods [verspoor2013nlp]. NLP encompasses a variety of tasks aimed at understanding, processing, and generating human language. Named entity recognition (NER), a prominent NLP task, focuses on identifying and classifying entities within the text, such as names of individuals, medical conditions, etc. For instance, in medical literature, NER can assist in extracting crucial information from documents. Text summarization in NLP is widely used for generating coherent summaries of lengthy texts. Sentiment analysis is a task that determines the emotional tone expressed in a given text, providing valuable insights for applications like social media monitoring or customer feedback analysis. Machine translation is a fundamental NLP task in breaking down language barriers by automatically translating text from one language to another. Question answering in NLP is directed at comprehending and responding to user queries, propelling advancements in virtual assistants and information retrieval.
2.2.1 Tokenization
@@
The first step in NLP is tokenization, which is the mechanism of splitting or fragmenting the sentences and words to their possible smallest morpheme, called a token. A morpheme is the smallest possible word after which it cannot be broken further [rai2021token]. One example of a word-level tokenization method is whitespace tokenization, which segments text based on whitespace characters. In many NLP applications, subword tokenization methods are preferred due to their effectiveness in handling out-of-vocabulary words. WordPiece [wu2016wordpiece] begins by treating each character as a token, creating an initial vocabulary. Employing a flexible and adaptive merging strategy, WordPiece considers any pair of adjacent characters or subword units that enhance the overall likelihood of the training data. This likelihood reflects the model’s probability of accurately representing the training data given its current state. In contrast, Byte-Pair Encoding (BPE) [sennrich2016bpe] shares similarities with WordPiece but adheres to a more deterministic merging strategy. In each iteration, BPE merges the most frequent pair of adjacent characters or subword units, progressing toward a predefined vocabulary size. Byte-level BPE [wang2020bbpe] operates at an even finer granularity, considering individual bytes rather than characters. Byte-level BPE extends the concept of subword tokenization to bytes, allowing it to capture more nuanced patterns at the byte level.
2.2.2 Token Embeddings
@@
Tokens are then often transformed into numerical vectors that capture semantic relationships between tokens, which are referred to as word or token embeddings. Word2Vec [mikolov2013wordvec] is a widely used word embedding technique that uses two models: Skip-Gram [mikolov2013wordvec] and Continuous Bag of Words (CBOW) [mikolov2013cbow]. In skip-gram, the model predicts context words given a target word, capturing semantic associations. Conversely, CBOW predicts the target word based on its context, emphasizing syntactic structures. In both models, the “context word” refers to words within a specified window around the target word, and the “target word” is the word for which predictions are made. Word2Vec is computationally efficient, making it well-suited for large datasets and general-purpose applications. Global Vectors (GloVe) [pennington2014glove] is a word embedding model that distinguishes itself by capturing global semantic relationships. It focuses on the entire corpus rather than local context windows. The model builds a co-occurrence matrix based on the global statistics of word pairs and then employs an objective function to generate word vectors that reflect the ratio of co-occurrence probabilities. GloVe uses an implicit skip-gram approach, capturing word relationships globally, making it ideal for tasks that demand a holistic understanding of word connections. FastText [bojanowski2017fasttext] is another word embedding that is particularly effective for handling out-of-vocabulary words and morphologically rich languages. It adopts a sub-word approach, breaking words into n-grams, and employs a skip-gram training method similar to Word2Vec [mikolov2013wordvec] to learn embeddings for these sub-word units. There are also word embeddings tailored to represent biomedical and clinical terms better. In tasks where the order of words is not essential, other feature extraction techniques can be effective, for example, bag-of-words (BoW) in text classification or term frequency-inverse document frequency (tf-idf) for information retrieval.
@@
In addition to general-purpose word embeddings, there are ones designed for biomedical and clinical terms. BioWordVec [zhang2019biowordvec] incorporates MeSH (Medical Subject Headings) terms along with text from PubMed abstracts and employs the fastText [bojanowski2017fasttext] algorithm to learn improved biomedical word embeddings. Another prominent approach is Cui2vec [beam2018cui2vec], which leverages diverse multi-modal data derived from medical publications and clinical notes. Cui2vec systematically maps medical terms onto a common Concept Unique Identifier (CUI) space, followed by the construction of a co-occurrence matrix. This matrix captures instances where different CUIs appear together, which is a foundation for generating word embeddings using techniques such as GloVe [pennington2014glove] or Word2Vec [mikolov2013wordvec]. In most cases, it is common to add positional encodings to capture the order of tokens in a sequence. Positional encoding vectors, often based on sinusoidal functions, systematically encode token positions, enriching embeddings with positional information for utilization in ML models tailored to specific NLP tasks [ahmed2023transformers].
2.2.3 Recurrent Neural Networks (RNNs)
@@
RNNs are widely employed for pattern detection in sequential data, encompassing diverse types such as genomic sequences, text, or numerical time series [schmidt2019rnn]. Operating on the principle of preserving a form of memory, RNNs incorporate a cyclic structure by looping the output of a specific layer back to the input, facilitating the prediction of subsequent layer outputs. This mechanism empowers RNNs to adeptly model sequential and temporal dependencies, capturing information from preceding time steps within hidden states. Despite their capacity to retain information from past inputs, RNNs encounter challenges in preserving long-term dependencies within input sequences due to the vanishing gradient problem. To address this, several RNN variants, including Long Short-Term Memory (LSTM) [hochreiter1997lstm] and Gated Recurrent Unit (GRU) [cho2014gru] networks, have been designed to enhance their ability to capture and utilize long-range dependencies in sequential data.
2.2.4 Transformers